OpenStudy (anonymous):

Could someone help me do this, Solve: (express the answer using interval notation.) l4x-7l (greater then or equal to) 5

OpenStudy (anonymous):

\[|4x-7| \ge 5 \] could someone help please?

OpenStudy (anonymous):

the trick with absolute values is to analyse them as two different equations 1) \[4x-7\le5\] \[4x \le12\] \[x \le 3\] 2) \[4x-7\ge-5\] \[4x\ge2\] \[x\ge0.5\] therefore x has to be between 0.5 and 3 therefore \[0.5 \le x \le 3\]

OpenStudy (anonymous):

Ohh, thanks a lot for your help, I appreciate it.